Find the equation to the straight line cutting of intercept (-5) and 6 from the axis

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
x intercept = -5 and y intercept = 6
 
hence equation of straight line is-
 
x/ –5  + y/6 = 1
 
 – 6x + 5y = 30
Vikas TU
4 years ago
Der student 
Equation of straight line having intercept a, and b is given by 
X/a + Y/b = 1 
so. The equation would be 
-X/5 + Y/6 =0 
-6x + 5y =30